Air is pushed into a soap bubble of radius r to double its radius. if the surface tension of the soap solution is S the work done in the process is
Options :
(a) 8 pi r^2 S
(b) 12 pi r^2 S
© 16 pi r^2 S
(d) 24 pi r^2 S

Grade:12

2 Answers

Manish Mohapatra
7 years ago
Your answer will b c I.e16pi r^2s.................it is because radius of 2 water bubbles ratio will b 16:9
Shivam Jain
7 years ago
Initial radius of bubble r1 = rFinal radius r2 = 2rSurface tension = SIncrease in surface area = ds = 2× 4π × *((2r)^2 – (r)^2) = 8π(3r^2) = 24πr^2 S Work done = S×ds=24πr^2